OpenStudy (anonymous):

1)3x-10 when x=5 2)3x-10 when x=6 3)3x-10 when x=-2 4)3x-10 when x=1.5

OpenStudy (anonymous):

1) when x = 5 then 3x5-10 = 15-10 =5 2)when x=6 then 3x6-10= 18-10 =8 3)when x= -2 then 3x(-2)-10 = -6-10 =-16 4)when x=1.5 then 3x1.5-10 = 4.5-10=5.5